OpenStudy (anonymous):

Find the vertex, y-intercept and x-intercepts of f(x)=-3x^2-6x+24. Please show how you did it. Thank you

OpenStudy (anonymous):

the vertex is (-b/2a, f(-b/2a)) = (6/-6, f(6/-6)) = (-1, f(-1)) = (-1, 27) y-intercept is for x = 0: (0, 24) x-intercept is for y = 0: -3(x^2 + 2x - 8) = 0, -3(x - 2)(x + 4) = 0, x = 2, -4 x-intercepts are (2, 0), (-4, 0)

OpenStudy (anonymous):

Thank you